# Collaborative Agents for Software Engineering

### v2

@@ 5 Experimental Result Analysis
In this Section, we discuss the performance of CodeAgent on four main experiments: vulnerability analysis (Section 5.1), inconsistency detection between commit and commit message (Section 5.2), format inconsistency detection (Section 5.2), and code revision (Section 5.3). We also discuss the difference in execution time of CodeAgent across different languages and conduct capabilities analysis between CodeAgent and other communicative agents in Appendix-Section D.
@@ 5.1 Vulnerability Analysis
As shown in Table 1, vulnerability analysis ❷ is one subtask of code reviews. Compared to ❶ and ❸, ❷ is a more complex code review subtask, covering more than 25 different aspects (please see the Appendix-Section F), including buffer overflows, sensitive data exposure, configuration errors, data leakage, etc. This domain necessitates deep technical expertise for accurate data annotation, thereby significantly increasing the time and money cost when labeling manually. Considering the low proportion of commits with vulnerabilities, this paper proposes a proactive verification method for data annotation.
Our approach is twofold. First, we utilize CodeAgent to process 3,545 pairs of commits, commit messages and original files across nine languages. These identified data points for potential vulnerabilities were then subjected to manual verification. Second, we employed tools like CodeBERT [12] and ChatGPT to do vulnerability binary detection in the same dataset. The results were then verified for their authenticity.
| Approach | CodeBERT | ChatGPT-3.5 | ChatGPT-4.0 | CodeAgent | CodeAgent |
| Find | 1,063 | 864 | 671 | 483 | 564 |
| Confirm | 212 | 317 | 345 | 449 | 413 |
| Rate | 19.94% | 36.69% | 51.42% | 92.96% | 73.23% |
| Rate | 5.98% | 8.94% | 9.73% | 12.67% | 11.65% |
@@
As delineated in Table 3, the deployment of CodeAgent successfully identified 483 potential vulnerabilities within a dataset of 3,545 items, with an impressive 449 of these ultimately confirmed as high-risk vulnerabilities, substantiated through a rigorous manual verification process exceeding 120 working hours111The verification process involved meticulous manual examination, extending beyond 120 working hours.. CodeBERT, a key pre-trained model for code-related tasks, with its parameters frozen for this experiment, initially identified 1,063 items as vulnerable, yet only 212 passed the stringent verification criteria. Similar trends were observed with ChatGPT-3.5 and ChatGPT-4.0, which confirmed 317 and 345 vulnerabilities out of 864 and 671 identified items, respectively. These outcomes are further quantified by the confirmation rates (Rate) of 19.94% for CodeBERT, 36.69% for ChatGPT-3.5, and 51.42% for ChatGPT-4.0, while CodeAgent demonstrated a remarkable Rate of 92.96%. Additionally, the analysis of confirmed vulnerabilities against all analyzed items (Rate) yielded 5.98%, 8.94%, 9.73%, and 12.67% for CodeBERT, ChatGPT-3.5, ChatGPT-4.0, and CodeAgent, respectively. Evidently, Table 3 not only highlights CodeAgent’s high precision in identifying vulnerable commits but also reveals the progressive improvement from ChatGPT-3.5 to ChatGPT-4.0, likely due to the latter’s capacity to handle longer input sequences, with token limits of 4,096 and 32,768, respectively. The integration of sophisticated algorithms like CoT and QA-Checker in CodeAgent has significantly enhanced its capabilities in vulnerability detection, surpassing the individual input-output efficiencies of ChatGPT and CodeBERT. Further details regarding the importance of the QA-checker can be seen in Appendix-Section C and Section L. Moreover, more experimental results in 9 languages are accessible in Appendix-Section I.
@@
In addition, the analysis of vulnerabilities identified by various models reveals interesting overlaps in their findings. CodeBERT confirmed 212 vulnerabilities, whereas ChatGPT-3.5, ChatGPT-4.0, and CodeAgent confirmed 317, 345, and 449 vulnerabilities, respectively. Notably, the intersection of vulnerabilities confirmed by CodeBERT and ChatGPT-3.5 is 169, indicating a substantial overlap in their findings. Similarly, the intersection between CodeBERT and ChatGPT-4.0 is 170, while a more significant overlap of 212 vulnerabilities is observed between ChatGPT-3.5 and ChatGPT-4.0. The combined intersection among CodeBERT, ChatGPT-3.5, and ChatGPT-4.0 is 137, underscoring the commonalities in vulnerabilities detected across these models. Furthermore, the intersections of vulnerabilities confirmed by CodeBERT, ChatGPT-3.5, and ChatGPT-4.0 with CodeAgent are 212, 317, and 334, respectively, highlighting the comprehensive coverage and detection capabilities of CodeAgent.
Ablation Study
As shown in Table 3, without QA-Checker, CodeAgent is less effective in finding vulnerable issues and reduces the hit rate (Rate and Rate) compared with the full version, indicating the importance of QA-Checker in our CodeAgent. More detailed information about the ablation study can be found in Appendix-Section L.
Moreover, CodeAgent versions 3.5 and 4.0 yielded consistent detection results, but they differ in the quality of explanation when doing the vulnerability analysis of the commit. More details about the difference between CodeAgent-3.5 and CodeAgent-4.0 are provided in Appendix-Section K.2.
@@ 5.2 Consistency and Format Detection
In this Section, we will discuss the performance of CodeAgent and baselines on metrics like the F1-Score and recall score of task ❶ and ❸. For ❶ and ❸, the dataset we have is shown in Table 2 and more detailed data information is shown in Figure 7 in Appendix.
Consistency Detection Between Commit and Commit Message
@@
Our comprehensive study, as illustrated in Table 4, assesses CodeAgent’s efficacy in detecting the consistency between commit and commit message, contrasting its performance with other prevalent methods like CodeBERT, ChatGPT-3.5, and ChatGPT-4.0. This evaluation specifically concentrates on merged and closed commits across nine languages, with a keen focus on crucial metrics such as Recall and F1-Score. Notably, CodeAgent exhibits a remarkable performance, outstripping other methods in both merged and closed scenarios. In terms of Recall, CodeAgent achieved an impressive 90.11% for merged commits and 87.15% for closed ones, marking a considerable average improvement of 5.62pp over the other models. Similarly, the F1-Score of CodeAgent stands at 93.89% for merged and 92.40% for closed commits, surpassing its counterparts with an average improvement of 4.00pp. More comparable details in different languages are shown in Appendix-Section. J.
| Merged | CodeBERT | ChatGPT-3.5 | ChatGPT-4.0 | CodeAgent | Imp (pp) |
| Recall | 63.64 | 80.08 | 84.27 | 90.11 | 5.84 |
| F1 | 75.00 | 87.20 | 90.12 | 93.89 | 3.77 |
| Closed | CodeBERT | ChatGPT-3.5 | ChatGPT-4.0 | CodeAgent | Imp (pp) |
| Recall | 64.80 | 79.05 | 81.75 | 87.15 | 5.40 |
| F1 | 77.20 | 87.35 | 89.10 | 92.40 | 3.30 |
| Average | CodeBERT | ChatGPT-3.5 | ChatGPT-4.0 | CodeAgent | Imp (pp) |
| Recall | 64.22 | 79.57 | 83.01 | 88.63 | 5.62 |
| F1 | 76.01 | 87.28 | 89.61 | 93.16 | 4.00 |
Format Consistency Detection Between Commit and Original File
@@
In our detailed evaluation of format consistency between commits and original files, CodeAgent’s performance was benchmarked against established models like CodeBERT and ChatGPT variants across nine different languages. This comparative analysis, presented in Table 5, was centered around pivotal metrics such as Recall and F1-Score. CodeAgent demonstrated a significant edge over the state-of-the-art, particularly in the merged category, with an impressive Recall of 89.34% and an F1-Score of 94.01%. These figures represent an average improvement of 10.81pp in Recall and 6.94pp in F1-Score over other models. In the closed category, CodeAgent continued to outperform, achieving a Recall of 89.57% and an F1-Score of 94.13%, surpassing its counterparts with an average improvement of 15.96pp in Recall and 9.94pp in F1-Score. The overall average performance of CodeAgent further accentuates its superiority, with a Recall of 89.46% and an F1-Score of 94.07%, marking an average improvement of 15.96pp in Recall and 10.45pp in F1-Score. These results underscore CodeAgent’s exceptional capability in accurately detecting format consistency between commits and their original files.
| Merged | CodeBERT | ChatGPT-3.5 | ChatGPT-4.0 | CodeAgent | Imp (pp) |
| Recall | 60.59 | 60.72 | 78.53 | 89.34 | 10.81 |
| F1 | 74.14 | 74.88 | 87.07 | 94.01 | 6.94 |
| Closed | CodeBERT | ChatGPT-3.5 | ChatGPT-4.0 | CodeAgent | Imp (pp) |
| Recall | 69.95 | 73.61 | 68.46 | 89.57 | 15.96 |
| F1 | 80.49 | 84.19 | 80.16 | 94.13 | 9.94 |
| Average | CodeBERT | ChatGPT-3.5 | ChatGPT-4.0 | CodeAgent | Imp (pp) |
| Recall | 65.27 | 67.17 | 73.50 | 89.46 | 15.96 |
| F1 | 77.32 | 79.54 | 83.62 | 94.07 | 10.45 |
@@ 5.3 Code Revision
In this Section, we evaluate the effectiveness of various approaches in bug fixing by comparing their Error Percentage (EP) performance. The methods under consideration include Trans-Review, AutoTransform, T5-Review, CodeBERT, GraphCodeBERT, CodeT5, and CodeAgent. As detailed in Table 6, these approaches exhibit varied performance across different datasets. Notably, CodeAgent demonstrates a remarkable performance, particularly in the T5-Review dataset, where it achieves the highest EP of 37.6%. This is a significant improvement over other methods, underscoring the effectiveness of CodeAgent in handling complex code revision tasks. Additionally, with an average EP of 31.6%, CodeAgent consistently outperforms its counterparts, positioning itself as a leading solution in automated code revision. The contribution of CodeAgent to the field of automated code revision is noteworthy. Its ability to excel in the T5-Review dataset, a challenging benchmark, indicates a sophisticated understanding and handling of nuanced bugs. Moreover, its overall average performance surpasses that of other state-of-the-art models, highlighting its robustness and reliability. These results suggest that CodeAgent outperforms the state-of-the-art approaches in fixing buggy codes on metric on average.
@@ 6 Related Work
Automating Code Review Activities Our work contributes to automating code review activities, focusing on detecting source code vulnerabilities and maintaining code consistency. Related studies include Hellendoorn et al. [15], who addressed code change anticipation, and Siow et al. [29], who introduced CORE for code modification semantics. Hong et al. [18] proposed COMMENTFINDER for comment suggestions, while Tufano et al. [35] and Li et al. [23] developed tools for code review automation using models like T5CR and CodeReviewer, respectively. Recently, Lu et al. [25] incorporated large language models for code review, enhancing fine-tuning techniques.
Collaborative AI Collaborative AI, involving AI systems working towards shared goals, has seen advancements in multi-agent LLMs [31; 28], focusing on collective thinking, conversation dataset curation [39; 21], and sociological phenomenon exploration [27]. Research by Akata et al. [1] and Cai et al. [5] further explores LLM cooperation and efficiency. However, there remains a gap in integrating these advancements with structured software engineering practices [21; 28], a challenge our approach addresses by incorporating advanced human processes in multi-agent systems. For a complete overview of related work, please refer to our Appendix-Section A.
@@ 7 Conclusion
This paper introduces CodeAgent, a novel multi-agent-based framework designed to automate code reviews, leveraging an innovative QA-Checker system to maintain focus and alignment with the review’s objectives. CodeAgent demonstrates promising capabilities in detecting vulnerabilities, ensuring consistency between code changes and commit messages, and improving uniformity of code style. Our CodeAgent outperforms existing state-of-the-art solutions. By considering the specific characteristics of the code review process and incorporating the human-like conversational framework, CodeAgent significantly enhances efficiency and accuracy. Finally, we believe that our paper opens a new avenue for future software development collaboration practice and research.
8 Impact Statement
The adoption of our multi-agent-based code review framework promises to revolutionize how software development teams ensure code quality. By automating and enhancing the code review process, developers can allocate more time to creative and value-added tasks, leading to increased productivity and a higher standard of software craftsmanship. We expect that the adaptive learning capabilities of our approach contribute to a sustainable and evolving solution, capable of addressing the dynamic challenges in modern software development and machine learning.
Appendix A Complete Related Work

### v4

@@ 4 Experimental Result Analysis
This section discusses the performance of CodeAgent in the four tasks considered for our experiments. In Appendix Section E, we provide further analyses: we discuss the difference in the execution time of CodeAgent in different languages and perform a capability analysis between CodeAgent and recent approaches.
@@ 4.1 Vulnerability Analysis
Compared to CA and FA, VA is a more complex code review subtask, covering more than 25 different aspects (please see the Appendix-Section G), including buffer overflows, sensitive data exposure, configuration errors, data leakage, etc. Vulnerability analysis being a costly, time-consuming, resource-intensive and sensitive activity, only a low proportion of commits are labeled. We therefore propose a proactive method for data annotion: we execute CodeAgent on the 3,545 samples (covering nine languages) and manual verify the identified cases to build a ground truth. Then, we applied CodeBERT Feng et al. (2020) and GPT on the dataset with the task of vulnerability binary prediction.
@@
As shown in Table 2, CodeAgent successfully identified 483 potential vulnerabilities within a data set of 3,545 samples, with an impressive 449 of these finally confirmed as high-risk vulnerabilities111The verification process involved a rigorous manual examination, extending beyond 120 working hours. Each sample being validated by at least 2 people: a researcher and an engineer. CodeBERT, a key pre-trained model for code-related tasks, with its parameters frozen for this experiment, initially identified 1,063 items as vulnerable, yet only 212 passed the stringent verification criteria. Similar trends were observed with GPT-3.5 and GPT-4.0, which confirmed 317 and 345 vulnerabilities out of 864 and 671 identified items, respectively. These outcomes are further quantified by the confirmation rates (Ratecr) of 19.94% for CodeBERT, 36.69% for GPT-3.5, and 51.42% for GPT-4.0, while CodeAgent demonstrated a remarkable Ratecr of 92.96%. Additionally, the analysis of confirmed vulnerabilities against all analyzed items (Rateca) yielded 5.98%, 8.94%, 9.73%, and 12.67% for CodeBERT, GPT-3.5, GPT-4.0, and CodeAgent, respectively. Evidently, Table 2 not only highlights CodeAgent’s high precision in identifying vulnerable commits but also reveals the progressive improvement from GPT-3.5 to GPT-4.0, likely due to the latter’s capacity to handle longer input sequences, with token limits of 4,096 and 32,768, respectively. The integration of sophisticated algorithms like CoT and QA-Checker in CodeAgent has significantly enhanced its capabilities in vulnerability detection, surpassing the individual input-output efficiencies of GPT and CodeBERT. Appendix-Sections D and M highlight further details regarding the importance of the QA-checker. Moreover, more experimental results in 9 languages are accessible in Appendix-Section J.
@@
In addition, the analysis of vulnerabilities identified by various models reveals interesting overlaps among the models. CodeBERT confirmed 212 vulnerabilities, whereas GPT-3.5, GPT-4.0, and CodeAgent confirmed 317, 345, and 449 vulnerabilities, respectively. Notably, the intersection of vulnerabilities confirmed by CodeBERT and GPT-3.5 is 169, indicating a substantial overlap in their findings. Similarly, the intersection between CodeBERT and GPT-4.0 is 170, while a larger overlap of 212 vulnerabilities is observed between GPT-3.5 and GPT-4.0. The combined intersection among CodeBERT, GPT-3.5, and GPT-4.0 is 137, underscoring the commonalities in vulnerabilities detected across these models. Further, the intersections of vulnerabilities confirmed by CodeBERT, GPT-3.5, and GPT-4.0 with CodeAgent are 212, 317, and 334, respectively, highlighting the comprehensive coverage and detection capabilities of CodeAgent.
Ablation Study.
As shown in Table 2, we conducted an ablation study to evaluate the effectiveness of the QA-Checker in CodeAgent. Specifically, we created a version of our tool without the QA-Checker, referred to as CodeAgent w/o. We then compared this version to the full version of CodeAgent that includes the QA-Checker. The results demonstrate that CodeAgent w/o is substantially less effective in identifying vulnerable issues, yielding lower hit rates (Ratecr and Rateca). This reduction in performance highlights the critical role of the QA-Checker in enhancing CodeAgent’s overall effectiveness. More detailed information about the ablation study can be found in Appendix-Section M.
@@ 4.2 Consistency and Format Detection
In this section, we will discuss the performance of CodeAgent and baselines on metrics like the F1-Score and recall score of task CA and FA. For CA and FA, the dataset we have is shown in Table 1 and more detailed data information is shown in Figure 7 in Appendix.
Code Change and Commit Message Consistency Detection.
@@
As illustrated in Table 3, we assess the efficacy of CodeAgent in detecting the consistency between code changes and commit messages, contrasting its performance with other prevalent methods like CodeBERT, GPT-3.5, and GPT-4.0. This evaluation specifically focuses on merged and closed commits in nine languages. In particular, CodeAgent exhibits remarkable performance, outperforming other methods in both merged and closed scenarios. In terms of Recall, CodeAgent achieved an impressive 90.11% for merged commits and 87.15% for closed ones, marking a considerable average improvement of 5.62% over the other models. Similarly, the F1-Score of CodeAgent stands at 93.89% for merged and 92.40% for closed commits, surpassing its counterparts with an average improvement of 3.79%. More comparable details in different languages are shown in Appendix-Section. K.
| Merged | CodeBERT | GPT-3.5 | GPT-4.0 | COT | ReAct | CodeAgent | Imp (pp) |
| Recall | 63.64 | 80.08 | 84.27 | 80.73 | 82.04 | 90.11 | 5.84 |
| F1 | 75.00 | 87.20 | 90.12 | 87.62 | 88.93 | 93.89 | 3.77 |
| Closed | CodeBERT | GPT-3.5 | GPT-4.0 | COT | ReAct | CodeAgent | Imp (pp) |
| Recall | 64.80 | 79.05 | 81.75 | 81.77 | 83.42 | 87.15 | 5.21 |
| F1 | 77.20 | 87.35 | 89.61 | 89.30 | 89.81 | 92.40 | 3.35 |
| Average | CodeBERT | GPT-3.5 | GPT-4.0 | COT | ReAct | CodeAgent | Imp (pp) |
| Recall | 64.22 | 79.57 | 83.01 | 81.25 | 82.73 | 88.63 | 5.62 |
| F1 | 76.01 | 87.28 | 89.61 | 88.46 | 89.37 | 93.16 | 3.79 |
Format Consistency Detection.
@@
In our detailed evaluation of format consistency between commits and original files, CodeAgent’s performance was benchmarked against established models like CodeBERT and GPT variants across nine different languages. This comparative analysis, presented in Table 4, was centered around pivotal metrics such as Recall and F1-Score. CodeAgent demonstrated a significant edge over the state-of-the-art, particularly in the merged category, with an impressive Recall of 89.34% and an F1-Score of 94.01%. These figures represent an average improvement of 10.81% in Recall and 6.94% in F1-Score over other models. In the closed category, CodeAgent continued to outperform, achieving a Recall of 89.57% and an F1-Score of 94.13%, surpassing its counterparts with an improvement of 15.56% in Recall and 9.94% in F1-Score. The overall average performance of CodeAgent further accentuates its superiority, with a Recall of 89.46% and an F1-Score of 94.07%, marking an average improvement of 13.39% in Recall and 10.45% in F1-Score. These results underscore CodeAgent’s exceptional capability in accurately detecting format consistency between commits and their original files.
| Merged | CodeBERT | GPT-3.5 | GPT-4.0 | COT | ReAct | CodeAgent | Imp (pp) |
| Recall | 60.59 | 60.72 | 78.53 | 70.39 | 71.21 | 89.34 | 10.81 |
| F1 | 74.14 | 74.88 | 87.07 | 80.69 | 82.18 | 94.01 | 6.94 |
| Closed | CodeBERT | GPT-3.5 | GPT-4.0 | COT | ReAct | CodeAgent | Imp (pp) |
| Recall | 69.95 | 73.61 | 68.46 | 73.39 | 74.01 | 89.57 | 15.56 |
| F1 | 80.49 | 84.19 | 80.16 | 83.65 | 83.90 | 94.13 | 9.94 |
| Average | CodeBERT | GPT-3.5 | GPT-4.0 | COT | ReAct | CodeAgent | Imp (pp) |
| Recall | 65.27 | 67.17 | 73.50 | 71.89 | 72.61 | 89.46 | 15.96 |
| F1 | 77.32 | 79.54 | 83.62 | 82.17 | 83.04 | 94.07 | 10.45 |
@@ 4.3 Code Revision
We evaluate the effectiveness of CodeAgent in revision suggestion (i.e., bug fixing) based on Edit Progress (EP) metric. We consider Trans-Review, AutoTransform, T5-Review, CodeBERT, GraphCodeBERT, CodeT5 as comparable state of the art. As detailed in Table 5, these approaches exhibit a varied performance across different datasets. In particular, CodeAgent shows remarkable performance in the T5-Review dataset, achieving the highest EP of 37.6%. This is a significant improvement over other methods, which underlines the effectiveness of CodeAgent in handling complex code revision tasks. Furthermore, with an average EP of 31.6%, CodeAgent consistently outperforms its counterparts, positioning itself as a leading solution in automated code revision. Its ability to excel in the T5-Review, a challenging benchmark data, indicates a strong capability to address complex bugs. In addition, its overall average performance surpasses other state-of-the-art models, highlighting its robustness and reliability.
@@ 5 Related Work
Automating Code Review Activities. Our work contributes to automating code review activities, focusing on detecting source code vulnerabilities and maintaining code consistency. Related studies include Hellendoorn et al. Hellendoorn et al. (2021), who addressed code change anticipation, and Siow et al. Siow et al. (2020), who introduced CORE for code modification semantics. Hong et al. Hong et al. (2022) proposed COMMENTFINDER for comment suggestions, while Tufano et al. Tufano et al. (2021) and Li et al. Li et al. (2022) developed tools for code review automation using models like T5CR and CodeReviewer, respectively. Recently, Lu et al. Lu et al. (2023) incorporated large language models for code review, enhancing fine-tuning techniques.
Collaborative AI. Collaborative AI, involving AI systems working towards shared goals, has seen advancements in multi-agent LLMs Talebirad and Nadiri (2023); Qian et al. (2023), focusing on collective thinking, conversation dataset curation Wei et al. (2023); Li et al. (2023a), and sociological phenomenon exploration Park et al. (2023). Research by Akata et al. Akata et al. (2023) and Cai et al. Cai et al. (2023) further explores LLM cooperation and efficiency. However, there remains a gap in integrating these advancements with structured software engineering practices Li et al. (2023a); Qian et al. (2023), a challenge our approach addresses by incorporating advanced human processes in multi-agent systems. For a complete overview of related work, please refer to our Appendix-Section B.
@@ 6 Conclusion
In this paper, we introduced CodeAgent, a novel multi-agent framework that automates code reviews. CodeAgent leverages its novel QA-Checker system to maintain focus on the review’s objectives and ensure alignment. Our experiments demonstrate CodeAgent’s effectiveness in detecting vulnerabilities, enforcing code-message consistency, and promoting uniform code style. Furthermore, CodeAgent outperforms existing state-of-the-art solutions in code revision suggestions. By incorporating human-like conversational elements and considering the specific characteristics of code review, CodeAgent significantly improves both efficiency and accuracy. We believe this work opens exciting new avenues for research and collaboration practices in software development.
Limitations
Firstly, the generalizability of the system across different software development environments or industries may require further validation and testing. While the system has shown promising results in the provided datasets, its applicability to other contexts remains uncertain without additional empirical evidence. This limitation suggests that the findings may not be fully transferable to all settings within the software development domain. Secondly, the baseline test used in the study might be insufficient. The current testing approach may not fully capture the system’s performance, particularly in edge cases or more complex scenarios. This could result in an overestimation of the system’s capabilities and an underestimation of its limitations. Further, more comprehensive testing is needed to establish a more robust baseline and to ensure that the system performs reliably across a wider range of conditions.
Ethics Statements
This study was conducted in compliance with ethical guidelines and standards for research. The research did not involve human participants, and therefore, did not require informed consent or ethical review from an institutional review board. All data used in this study were publicly available, and no personal or sensitive information was accessed or processed. The development and evaluation of the CodeAgent system were performed with a focus on transparency, reproducibility, and the potential positive impact on the software development community.
Appendix A Details of QA-Checker Algorithm
Lemma A.1.
Let denote the quality assessment function of the QA-Checker for the question-answer pair in a conversation at the -th iteration. Assume is twice differentiable and its Hessian matrix is positive definite. If the QA-Checker modifies the question to by attaching an additional instruction , and this leads to a refined answer , then the sequence converges to an optimal question-answer pair , under specific regularity conditions.
Proof.
The QA-Checker refines the question and answers using the rule:
Substituting the update rule and rearranging, we get:
In practical terms, this lemma and its proof underpin the QA-Checker’s ability to refine answers iteratively. The QA-Checker assesses the quality of each answer concerning the posed question, employing advanced optimization techniques that are modeled by the modified Newton-Raphson method to enhance answer quality. This framework ensures that, with each iteration, the system moves closer to the optimal answer, leveraging both first and second-order derivatives for efficient and effective learning.
Further Discussion
The QA-Checker computes at each iteration and compares it to a predefined quality threshold . If , the QA-Checker generates an additional instruction to refine the question to , prompting the agents to generate an improved answer .
First, we assume that the quality assessment function is twice differentiable with respect to the question . This assumption is reasonable given the smooth nature of the component functions (relevance, specificity, and coherence) and the use of continuous word embeddings. Next, we apply the second-order Taylor approximation to around the point :
where , is the Hessian matrix of evaluated at , and is the remainder term.
Assuming that the remainder term is negligible and that the Hessian matrix is positive definite, we can approximate the optimal step as:
Substituting this approximation into the Taylor expansion and using the fact that (where is the learning rate), we obtain:
The assumptions of twice differentiability, negligible remainder term, and positive definite Hessian matrix provide a more solid foundation for the approximation in Lemma 3.1. For sufficiently small , this approximation suggests an increase in , implying convergence to an optimal question-answer pair as . The convergence relies on the positive definiteness of and the appropriate choice of , ensuring each iteration moves towards an improved quality of the question-answer pair.
The quality assessment function used by the QA-Checker is defined as:
where:
and represent the question and answer at the -th iteration of the conversation.
measures how well the answer addresses the key points and intent of the question , computed as:
where and are vector representations of and .
assesses how specific and detailed the answer is, calculated as:
where is the set of substantive content words in , is a measure of how technical or domain-specific the term is, and is the total number of words in .
evaluates the logical flow and structural coherence of the answer , computed as:
where is the density of discourse connectives in , measures the consistency of coreference chains in , and assesses how well follows the expected structural patterns for the given question type.
, , and are non-negative weights that sum to 1, with .
Appendix B Complete Related Work